# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ 4 Delayed Full-Information Feedback
Lemma 4.2 with $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ in (2) and Cauchy–Schwarz yields $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})$, where $S_{1:T}=\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$. Hence, from Assumption 4.1, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})-aS_{1:T}=O(D^{2/3}T^{1/3}),$ where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. ∎
We can also prove the following high-probability bound; see Section E.3 for the proof:
@@
For any $\delta\in(0,1)$, with probability at least $1-\delta$, the above algorithm achieves $\mathcal{R}_{T}=O(\log({1}/{\delta})+D^{2/3}T^{1/3}).$
5 Delayed Bandit Feedback
Given the results so far, it is natural to explore online structured prediction with delayed bandit feedback. We construct algorithms for this setup by combining the theoretical developments from Section 3 and Section 4. This section assumes Assumption 3.2, as in Section 3.
5.1 $O(\sqrt{DKT})$ Regret Algorithm
We use RDUE with $q=B\sqrt{DK/T}$ as the decoding function (assuming $T\geq DB^{2}K$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (4), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$. This algorithm attains the following bound:
Theorem 5.1.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{DKT}).$
The proof can be found in Section F.2. Due to the introduction of the delay, the regret bound worsens by a factor of $\sqrt{D}$ compared to the non-delayed case, which is natural when considering analyses of adversarial cases in delayed feedback (Joulani et al., 2013; Zimmert & Seldin, 2020; Manwani & Agarwal, 2023).
5.2 $O(D^{1/3}T^{2/3})$ Regret Algorithm
Here, we make the same assumptions on the target loss function as in Section 3.4. We provide an algorithm that improves the dependence on $K$ from Section 5.1. We use RDUE with $q=\left(\frac{\omega B^{2}C_{x}^{2}D}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq\omega B^{2}C_{x}^{2}D$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (5), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$. This algorithm achieves the following bound:
@@ 5.2 $O(D^{1/3}T^{2/3})$ Regret Algorithm
The proof can be found in Section F.3. Due to the introduction of the delay, the regret bound worsens by a factor of $D^{1/3}$ compared to the non-delayed bandit feedback case.
6 Experiments
Figure 1: A box plot of error rate of the MNIST experiment for multiclass classification with bandit feedback.
This section presents numerical experiment results for online multiclass classification under bandit feedback. We compare three algorithms: Gaptron (Van der Hoeven, 2020) with logistic loss and hinge loss as surrogate losses, Gappletron (Van der Hoeven et al., 2021) with logistic loss as the surrogate loss, and our proposed algorithm from Section 5.1. The parameters for each algorithm are set based on their theoretical values. We use the MNIST dataset (LeCun et al., 2010), a dataset of digit images. The diameter $B$ of $\mathcal{W}$ is fixed at $10$. We repeated the experiment for 20 times, and the boxplot of the obtained misclassification rates is summarized in Figure 1. From Figure 1, we observe that our method achieves the lowest misclassification rate. Despite not being specialized for multiclass classification, our approach outperforms existing algorithms designed for multiclass classification on real data with $K=10$. Further experiments can be found in Appendix G, and related discussions are provided in Appendix C.
@@ 7 Conclusion
In this study, we developed multiple algorithms for online structured prediction under bandit and delayed feedback and analyzed their surrogate regret. Among several contributions, a noteworthy one is the algorithm for bandit feedback whose surrogate regret bound does not explicitly depend on the label set size $K$, which is achieved by leveraging the pseudo-inverse matrix estimator. Several interesting research directions remain. One is to investigate surrogate regret lower bounds under bandit feedback and delayed feedback. In the existing $\Omega(\sqrt{T})$ lower bound for bandit feedback, the optimal dependence on $d$ is unexplored. Investigating the tightness of the $O(D^{2/3}T^{1/3})$ upper bound for delayed and full information feedback is also an important question. Another key direction is extending our results to settings where the delay time is not constant in delayed feedback, which is crucial for some practical applications.
@@ Acknowledgments
TT is supported by JST ACT-X Grant Number JPMJAX210E and JSPS KAKENHI Grant Number JP24K23852, SS is supported by supported by JST ERATO Grant Number JPMJER1903, and KY is supported by JSPS KAKENHI Grant Number JP24H00703.
@@ Appendix A Notation
$C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$ diameter of input space $\mathcal{X}$
$C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ diameter of output space $\mathcal{Y}$
@@
$L:\mathcal{Y}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ target loss function
@@
$S_{\Omega}:\mathbb{R}^{n}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ Fenchel–Young loss generated by $\Omega$
@@
$L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$ upper bound of the gradient
@@ Appendix B Additional Related Work
We discuss additional related work that could not be included above.
@@
Before the introduction of the Fenchel–Young loss framework, Niculae et al. (2018) proposed SparseMAP, which used the squared $\ell_{2}$-norm regularization. The Fenchel–Young loss, described in Section 2.2, is built upon the idea of SparseMAP. The Structure Encoding Loss Function (SELF) was introduced by Ciliberto et al. (2016, 2020) to analyze the relationship between surrogate and target losses, a concept known as Fisher consistency. For more extensive literature, we refer the reader to Sakaue et al. (2024, Appendix A).
@@
In the full information setup, PERCEPTRON is one of the most representative algorithms for binary classification (Rosenblatt, 1958), and the multiclass setting has also been extensively studied (Crammer & Singer, 2003; Fink et al., 2006). Online logistic regression is another relevant research stream, with Foster et al. (2018) being a particularly representative study. The study of the bandit setup was initiated by Kakade et al. (2008), and it has since been extensively explored in subsequent research (Hazan & Kale, 2011; Beygelzimer et al., 2017; Foster et al., 2018). However, to the best of our knowledge, no prior work has addressed general structured prediction under bandit feedback. A most related study is the work by Gentile & Orabona (2014b), which investigated online multilabel classification and ranking. However, their setting assumes access to feedback of the form $\{\mathds{1}[\bm{y}_{t,i}\neq\widehat{\bm{y}}_{t,i}]\}_{i}$, which is more informative than bandit feedback and differs from our setup. Van der Hoeven (2020) explicitly introduced the surrogate regret in the context of online multiclass classification. This study has been extended to the setting where observations are determined by a directed graph (Van der Hoeven et al., 2021) and to structured prediction scenarios (Sakaue et al., 2024). For a more extensive overview of the literature on online classification, we refer the reader to Van der Hoeven (2020).
@@
The study of delayed feedback began with Weinberger & Ordentlich (2002). Since then, it has been extensively explored in various online learning settings, primarily in the full information setup of online convex optimization (Mesterharm, 2005; Joulani et al., 2013, 2016; Flaspohler et al., 2021). Algorithms for delayed bandit feedback have been studied mainly in the context of multi-armed bandits and their variants (Cesa-Bianchi et al., 2016; Zimmert & Seldin, 2020; Ito et al., 2020; Masoudian et al., 2022; Van der Hoeven et al., 2023). In the context of online classification, research considering delay is scarce; the only work is Manwani & Agarwal (2023) to our knowledge.
@@ Appendix C Discussion on the Difference in Surrogate Loss Functions
As in (1), the surrogate regret $\mathcal{R}_{T}$ is defined by $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$, which means the choice of the surrogate loss $S$ affects the bound on the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. Van der Hoeven et al. (2021, Theorem 1), which applies to a more general setup than bandit feedback, implies $\mathcal{R}_{T}=O(K\sqrt{T})$ for the bandit setup with $S$ being a logistic loss defined with the base-$K$ logarithm. On the other hand, our bound of $\mathcal{R}_{T}=O(\sqrt{KT})$ applies to the logistic loss $S$ defined with the base-$2$ logarithm. As a result, while our bound on $\mathcal{R}_{T}$ is better, the $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$ term can be worse; this is why we cannot directly compare our $O(\sqrt{KT})$ bound with the $O(K\sqrt{T})$ bound in Van der Hoeven et al. (2021, Theorem 1). We may use the decoding procedure in Van der Hoeven et al. (2021), instead of RDUE, to recover their bound that applies to the base-$K$ logistic loss. It should be noted that their method is specific to multiclass classification; naively extending their method to structured prediction formulated as $|\mathcal{Y}|$-class classification results in the undesirable dependence on $K=|\mathcal{Y}|$, as is also discussed in Sakaue et al. (2024). By contrast, our pseudo-inverse estimator, combined with RDUE, can rule out the explicit dependence on $K$, at the cost of the increase from $\sqrt{T}$ to $T^{2/3}$.
Appendix D Omitted Details of Section 3
@@ D.1 Concentration inequality
To prove high probability regret bounds, we use the following concentration inequality.
Lemma D.1 (Bernstein’s inequality, e.g., Cesa-Bianchi & Lugosi 2006, Lemma A.8).
@@
D.2 Proof of Theorem 3.6
Here, we provide the proof of Theorem 3.6. Hereafter, we let $S_{\max}=\max_{\bm{W}\in\mathcal{W}}S_{t}(\bm{W})$ and $\widehat{S}_{t}(\bm{W})=v_{t}S_{t}(\bm{W})=\frac{\mathds{1}[\bm{y}_{t}=\widehat{\bm{y}}_{t}]}{p_{t}(\widehat{\bm{y}}_{t})}S_{t}(\bm{W})$. The following theorem is the formal version of Theorem 3.6:
Theorem D.2 (Formal version of Theorem 3.6).
Consider the bandit and non-delayed setup. Let
@@
$\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right)$ $\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}=\sqrt{2}B\sqrt{\sum_{t=1}^{T}v_{t}^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}$
@@
Hence from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime\prime}\in(0,1)$, with probability at least $1-\delta^{\prime\prime}$ we have
@@
$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT\leq\sum_{t=1}^{T}v_{t}S_{t}(\bm{U})+\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-v_{t}S_{t}(\bm{W}_{t})\right)+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
$=\sum_{t=1}^{T}v_{t}\underbrace{\left(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})\right)}_{=r_{t}}-(1-\xi)\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
@@
where we recall that $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$. Rearranging the last inequality and using $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$ give
@@
$\mathcal{R}_{T}$ $\leq\sqrt{2\log(2/\delta)\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT+\sqrt{6\log(2/\delta)\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}$
$\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}$
@@
where the first inequality follows from (7) and $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$, and the second inequality follows from Assumption 3.2, the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Substituting $q=\sqrt{\mathcal{C}/2}$ and $\delta^{\prime}=\delta/2$ and in the last inequality, we obtain
@@
Here, we provide the proof of Theorem 3.8. We recall that $\bm{P}_{t}=\mathbb{E}_{t}\left[{\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}}\right]$. We then estimate $\bm{y}_{t}$ by $\widetilde{\bm{y}}_{t}={\bm{V}}^{-1}\bm{P}^{+}_{t}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle$ and $\bm{G}_{t}$ by $\widetilde{\bm{G}}_{t}\coloneqq(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}$ under Assumption 3.7. This $\widetilde{\bm{G}}_{t}$ satisfies $\mathbb{E}_{t}\left[{\widetilde{\bm{G}}_{t}}\right]=\bm{G}_{t}.$ To prove Theorem 3.8, we will upper bound $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$. To do so, we begin by proving the following lemma:
@@
Using this lemma we prove a property of $\bm{P}_{t}$ and an upper bound of $\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}\right)}\right]$. In what follows, we use $\lambda_{\min}(\bm{A})$ to denote the minimum eigenvalue of a matrix $\bm{A}$.
@@
Suppose that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where we recall $\mu$ is the uniform distribution over $\mathcal{Y}$. Then, we have
@@
where the first inequality follows from $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})\leq L_{t}(\widehat{\bm{y}}_{t})\leq 1$ since $\bm{b}\geq 0$ and $c(\cdot)$ is non-negative and the last equality follows from $\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}=\bm{P}^{+}_{t}$. Hence,

### v2

@@ 4 Delayed full-information feedback
From Lemma 4.2 with $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ in (1) and Cauchy–Schwarz ineqaulity, we have $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}=O(D\sqrt{S_{1:T}})$, where $S_{1:T}=\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$. Thus, Assumption 4.1 implies $\mathbb{E}[\mathcal{R}_{T}]\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}=O(D\sqrt{S_{1:T}})-aS_{1:T}=O(D^{2})$, where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. ∎
We can also prove a high-probability surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\}+\log({1}/{\delta}))$, which holds with probability at least $1-\delta$. See Section E.3 for the proof.
5 Delayed bandit feedback
Given the results so far, it is natural to explore OSP with delayed bandit feedback. We construct algorithms for this setting by combining the theoretical developments from Sections 3 and 4. This section assumes Assumption 3.2, as in Section 3.
5.1 Algorithm for bandit delayed feedback with $O(\sqrt{(K+D)T})$ regret
We adopt RDUE with $q=B\sqrt{K/T}$ for decoding (assuming $T\geq B^{2}K$), the inverse-weighted gradient estimator $\widehat{\bm{G}}_{t}$, and ODAFTRL with AdaHedgeD as $\mathsf{ALG}$. Then, the following bound holds:
@@ 5.1 Algorithm for bandit delayed feedback with $O(\sqrt{(K+D)T})$ regret
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{(K+D)T}).$
The proof can be found in Section F.2. This upper bound incurs an additional $O(\sqrt{DT})$ compared to the non-delayed case. Whether this surrogate regret upper bound is tight or not remains open, as it is not easy to establish a corresponding lower bound. As mentioned above, an $\Omega(\sqrt{T})$ surrogate regret lower bound exists in the graph feedback setting [39]. However, no such lower bound is known for the bandit setting, and constructing lower bounds that take the delay into account would be more difficult.
5.2 Algorithm for bandit delayed feedback with $O(D^{1/3}T^{2/3})$ regret
We provide an algorithm that improves the dependence on $K$ from Section 5.1. We make the same assumptions on the target loss function as Section 3.4. We use RDUE with $q=\left({\omega B^{2}C_{x}^{2}D}/{T}\right)^{1/3}$ for decoding (assuming $T\geq\omega B^{2}C_{x}^{2}D$), the pseudo-inverse matrix estimator $\widetilde{\bm{G}}_{t}$, and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$. Then, the following bound holds:
@@ 5.2 Algorithm for bandit delayed feedback with $O(D^{1/3}T^{2/3})$ regret
The proof can be found in Section F.3. Due to the presence of the delay, the surrogate regret bound worsens by a factor of $D^{1/3}$ compared to the non-delayed bandit setting. Additionally, we present algorithms for the variable-delay setting, which we defer to Appendix G due to space limitation.
@@ 6 Conclusion
We have developed multiple algorithms for OSP under bandit and delayed feedback and analyzed their surrogate regret. Among these contributions, of particular note is the algorithm for bandit feedback whose surrogate regret bound does not explicitly depend on the output set size $K$, achieved by leveraging the pseudo-inverse matrix estimator.
Let us discuss limitations regarding tightness. As for the bandit setting, the lower bound of $\Omega(\sqrt{T})$ in the graph-feedback setting [39] suggests that a similar lower bound likely exists, but this has not yet been proven, leaving the tightness of our upper bounds open. In the delayed full-information setting, our $O(\min\{D^{2},D^{2/3}T^{1/3}\})$ upper bound is independent of $T$, and its gap to the lower bound, if any, is considered small; however, it remains open. In the delayed bandit setting, although the dependence on $D$ is less critical than that on $T$, the optimal dependence on $D$ is to be explored in future work.
@@ Acknowledgments
TT is supported by JST ACT-X Grant Number JPMJAX210E and JSPS KAKENHI Grant Number JP24K23852, SS was supported by JST ERATO Grant Number JPMJER1903, and KY is supported by JSPS KAKENHI Grant Number JP24H00703.
@@ Appendix A Notation
$C_{x}=\max_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}$ maximum norm of input vectors in $\mathcal{X}$
$C_{y}$ the maximum of the largest Euclidean norm of vectors in $\operatorname{conv}(\mathcal{Y})$ or the diameter of $\operatorname{conv}(\mathcal{Y})$
@@
$L\colon\mathcal{Y}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ target loss function
@@
$S_{\Omega}\colon\mathbb{R}^{n}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ Fenchel–Young loss generated by $\Omega$
@@
${\widehat{\bm{G}}_{t}}$ inverse weighted estimator
${\widetilde{\bm{G}}_{t}}$ pseudo-inverse matrix estimator
@@
$D\in\mathbb{N}$ fixed delay time
$\tau_{t}$ variable delay time
$\tau_{\ast}=\max_{t}\tau_{t}$ maximum value of delay time
$\rho(t)$ the time step of the $t$th feedback from SOLID to BASE
$\widetilde{\tau}_{t}=t-1-\sum_{s=1}^{\rho(t)-1}\mathds{1}[s+\tau_{s}<\rho(t)]$ the number of feedback from SOLID to BASE pending during the $t$th feedback
@@ Appendix B Additional related work
We discuss additional related work that could not be included in the main text.
@@
Before the development of the Fenchel–Young loss framework, Niculae et al. [30] proposed SparseMAP, which used the squared $\ell_{2}$-norm regularization. The Fenchel–Young loss, described in Section 2.2, is built upon the idea of SparseMAP. The Structure Encoding Loss Function (SELF) was introduced by Ciliberto et al. [10, 11] to analyze the relationship between surrogate and target losses, a concept known as Fisher consistency. For a more extensive literature review, we refer the reader to Sakaue et al. [35, Appendix A].
@@
In the full-information setting, PERCEPTRON is one of the most representative algorithms for binary classification [34], and the multiclass setting has also been extensively studied [13, 16]. Online logistic regression is another relevant research stream, with Foster et al. [18] being a particularly representative study. The study of the bandit setting was initiated by Kakade et al. [24], and it has since been extensively explored in subsequent research [20, 4, 18]. However, to the best of our knowledge, no prior work has addressed general structured prediction under bandit feedback. One of the most relevant studies is the work by Gentile and Orabona [19], who investigated online multilabel classification and ranking. However, their setting assumes access to feedback of the form $\{\mathds{1}[\bm{y}_{t,i}\neq\widehat{\bm{y}}_{t,i}]\}_{i}$, which is more informative than bandit feedback and differs from our setting. van der Hoeven [38] introduced the surrogate regret in the context of online multiclass classification. This study has been extended to the setting where observations are determined by a directed graph [39] and to structured prediction [35]. For a more extensive overview of the literature on online classification, we refer the reader to van der Hoeven [38].
@@
The study of delayed feedback was initiated by Weinberger and Ordentlich [41]. Since then, it has been extensively explored in various online learning settings, primarily in the full-information setting of online convex optimization [29, 22, 23, 17]. Algorithms for delayed bandit feedback have been studied mainly in the context of multi-armed bandits and their variants [9, 42, 21, 28, 40]. In online classification, research considering delay is scarce; the only work is that of Manwani and Agarwal [27] to our knowledge. There are several differences between their work and ours. Among them, a key distinction is that their study focuses on multiclass classification, whereas we address the more general OSP.
@@ Appendix C Discussion on the difference in surrogate loss functions
As in Section 1, the surrogate regret, $\mathcal{R}_{T}$, is defined by $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$, which means the choice of the surrogate loss function, $S$, affects the bound on the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. van der Hoeven et al. [39, Theorem 1], which applies to a more general setting than bandit feedback, implies $\mathcal{R}_{T}=O(K\sqrt{T})$ for the bandit setting with $S$ being a logistic loss defined with the base-$K$ logarithm. On the other hand, our bound of $\mathcal{R}_{T}=O(\sqrt{KT})$ applies to the logistic loss $S$ defined with the base-$2$ logarithm. As a result, while our bound on $\mathcal{R}_{T}$ is better, the $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$ term can be worse; this is why we cannot directly compare our $O(\sqrt{KT})$ bound with the $O(K\sqrt{T})$ bound in van der Hoeven et al. [39, Theorem 1]. We may use the decoding procedure in van der Hoeven et al. [39], instead of RDUE, to recover their bound that applies to the base-$K$ logistic loss. It should be noted that their method is specific to multiclass classification; naively extending their method to structured prediction formulated as $|\mathcal{Y}|$-class classification results in the undesirable dependence on $K=|\mathcal{Y}|$, as is also discussed in Sakaue et al. [35]. By contrast, our pseudo-inverse estimator, combined with RDUE, can rule out the explicit dependence on $K$, at the cost of the increase from $\sqrt{T}$ to $T^{2/3}$.
Appendix D Details omitted from Section 3
@@ D.1 Concentration inequality
To prove high probability regret bounds, we will use the following concentration inequality.
Lemma D.1 (Bernstein’s inequality, e.g., [7, Lemma A.8]).
@@
D.2 Proof of high probability bound
Here, we provide the proof of a high probability bound. Hereafter, we let $S_{\max}=\max_{\bm{W}\in\mathcal{W}}S_{t}(\bm{W})$ and $\widehat{S}_{t}(\bm{W})=v_{t}S_{t}(\bm{W})=\frac{\mathds{1}[\bm{y}_{t}=\widehat{\bm{y}}_{t}]}{p_{t}(\widehat{\bm{y}}_{t})}S_{t}(\bm{W})$. The following theorem is the formal version of the high probability bound under the bandit feedback:
Theorem D.2.
Consider the bandit and non-delayed setting. Let
@@ D.2 Proof of high probability bound
$\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right)$ $\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}=\sqrt{2}B\sqrt{\sum_{t=1}^{T}v_{t}^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}$
@@
Hence, from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime\prime}\in(0,1)$, with probability at least $1-\delta^{\prime\prime}$ we have
@@
$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT$
$\leq\sum_{t=1}^{T}v_{t}S_{t}(\bm{U})+\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-v_{t}S_{t}(\bm{W}_{t})\right)+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
$=\sum_{t=1}^{T}v_{t}\underbrace{\left(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})\right)}_{=r_{t}}-(1-\xi)\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})$
$\qquad+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
@@
where we recall that $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$. Rearranging the last inequality and using $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$ yield
@@
$\mathcal{R}_{T}$ $\leq\sqrt{2\log(2/\delta)\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT$
$\qquad+\sqrt{6\log(2/\delta)\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)$
$\qquad+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}$
@@
where the first inequality follows from (4) and $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$, and the second inequality follows from Assumption 3.2, the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Substituting $q=\sqrt{\mathcal{C}/T}$ and $\delta^{\prime}=\delta/2$ in the last inequality, we obtain
@@
Here, we provide the formal version and the proof of Theorem 3.7.
Theorem D.4 (Formal version of Theorem 3.7).
The algorithm in Section 3.4 achieves
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\omega^{1/3}\left(BC_{x}T\right)^{2/3}.$
We recall that $\bm{P}_{t}=\mathbb{E}_{t}\left[{\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}}\right]$. We then estimate $\bm{y}_{t}$ by $\widetilde{\bm{y}}_{t}={\bm{V}}^{-1}\bm{P}^{+}_{t}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle$ and $\bm{G}_{t}$ by $\widetilde{\bm{G}}_{t}\coloneqq(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}$ under Assumption 3.6. This $\widetilde{\bm{G}}_{t}$ satisfies $\mathbb{E}_{t}\left[{\widetilde{\bm{G}}_{t}}\right]=\bm{G}_{t}.$ To prove Theorem 3.7, we will upper bound $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$. To do so, we begin by proving the following lemma:
@@
Using this lemma, we prove a property of $\bm{P}_{t}$ and an upper bound of $\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}\right)}\right]$. In what follows, we use $\lambda_{\min}(\bm{A})$ to denote the minimum eigenvalue of a matrix $\bm{A}$.
@@
Suppose that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where we recall that $\mu$ is the uniform distribution over $\mathcal{Y}$. Then, we have
@@
where the first inequality follows from $-1\leq\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle\leq 1$ and the last equality follows from $\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}=\bm{P}^{+}_{t}$. The right-hand side is bounded as follows: